In a new system of units energy, density and power are taken as fundamental units, then the dimensional formula of momentum will be ---------?

Grade:11

2 Answers

Arun
5 years ago
\implies x=\dfrac{7}{5}, y=\dfrac{1}{5}, z=\dfrac{-3}{5}[Momentum]=[MLT^{-1}]=[Energy]^x[Density]^y[Power]^zPower = [ML^2 T^-3]Dear studentEnergy]=[ML^2T^{-2}]Density = [M L^-3]
 
Hope it helps. In case of any query please feel free to aks
Khimraj
5 years ago
Energy E = [ML2T-2]
Density D = [ML-3]
Power P = [ML2T-3]
Momentum = [MLT-1] = ExDyPz
SO x = 7/5, y = 1/5, z = -3/5
Then 
Momentum = [E7/5D1/5P-3/5].
Hope it clears.